OpenStudy (anonymous):

eight times the sum of a number and 6 equals 7

TheSmartOne (thesmartone):

sum of a number and 6 is x+6 so 8(x+6)=7 8x+48=7 8x=-41 x=-41/8

OpenStudy (anonymous):

yes lol you corrected it.

OpenStudy (immanuelv):

retiring from this site... giving too many wrong answers omg

OpenStudy (dan815):

lool